OpenStudy (anonymous):

a pair of shoes has an original price of $80 and they are on sale for $48, what is the percent of discount?

OpenStudy (anonymous):

http://www.mathgoodies.com/lessons/percent/old/solutions_unit10.html maybe it help[s

OpenStudy (stamp):

What I would do is say 80 times what percent p/100 = 48 So, solve 80(p / 100) = 48

OpenStudy (stamp):

80p/100 = 48 80p = 4800 p = 60 So $48 is 60% of $80, meaning the shoes were discounted 40%.

OpenStudy (stamp):

This is my method, it may differ from the way you are being taught in class, so if you do not understand, I apologize and you will require somebody else's assistance.

OpenStudy (anonymous):

well